# The Effect of Personalization in FedProx: A Fine-grained Analysis on Statistical Accuracy and Communication Efficiency

### v3

@@ C.1 Additional Experiment Details
Evaluation Criterion. In terms of evaluation, for the synthetic dataset, we track the ground-truth global model ${{\bm{w}}_{\star}^{(g)}}$ and local model ${{\bm{w}}_{\star}^{(i)}}$. We compute the global statistical error as $\|\widetilde{\bm{w}}^{(g)}-{{\bm{w}}_{\star}^{(g)}}\|^{2}$ and the local statistical error as $\|\widetilde{\bm{w}}^{(i)}-{{\bm{w}}_{\star}^{(i)}}\|^{2}$. To capture the total error at the $t$-th iteration, we report both the global total error $\|\bm{w}_{t}^{(g)}-{{\bm{w}}_{\star}^{(g)}}\|^{2}$ and the local total error $\|{\bm{w}}_{t,0}^{(i)}-{{\bm{w}}_{\star}^{(i)}}\|^{2}$, which combine both optimization and statistical errors. For the MNIST dataset, we report test accuracy on the MNIST testing set as a measure of the statistical error, and training loss for each client’s training set as total error at the $t$-th iteration.
C.2 Additional Experiment Results
Figure 3: Performance of the FedProx algorithm with an adaptive choice of $\lambda$ (from Theorem 1) under varying levels of statistical heterogeneity.
In Figure 3, statistical error of the FedProx algorithm with an adaptive choice of $\lambda$ under varying levels of statistical heterogeneity. With the adaptive choice of $\lambda$ as defined in Theorem 1, FedProx consistently achieves the lowest statistical error across most scenarios, outperforming both GlobalTrain and LocalTrain baselines. This is particularly evident under moderate and high levels of statistical heterogeneity. However, as $R$ approaches very small values, the global baseline performs slightly better, which aligns with the slower rate of FedProx with small $R$ described in our theoretical results. Overall, these results validate our choice of adaptive $\lambda$ and confirm the statistical rate provided by our analysis.
Figure 4: (a) Testing accuracy of different methods (FedProx, LocalTrain, GlobalTrain) versus statistical heterogeneity levels on the MNIST dataset (logistic regression). (b) Training loss versus communication rounds for the same methods.
In Figure 4(a), we compare FedProx, LocalTrain, and GlobalTrain on the MNIST dataset by plotting testing accuracy against varying levels of statistical heterogeneity. As heterogeneity increases, each client has access to fewer classes, resulting in fewer classification tasks. This change simplifies the problem, which explains why all methods show improved evaluation accuracy as the number of classes decreases. However, this does not diminish the insights from the relative comparison between methods. GlobalTrain and FedProx with larger $\lambda$ perform better at lower heterogeneity levels, while LocalTrain and FedProx with smaller $\lambda$ perform better as heterogeneity increases.
In Figure 4(b), we track testing loss as a function of communication rounds. Across all sub-figures, FedProx with smaller $\lambda$ (higher personalization) demonstrates faster convergence, especially in the early stages of training. The final training loss achieved by different methods also aligns with the results shown in Figure 4(a) and Figure 1.
Figure 5: Performance of FedProx with 100 clients (top row) and 1,000 clients (bottom row) with 10% client sampling under different levels of statistical heterogeneity ($R$).
In Figure 5, the top row shows results for 100 clients, and the bottom row shows results for 1,000 clients, and in both results we adopt a client sampling of 10% in each communication round. This tests the scalability and generalizability of the our results under typical federated learning setups. Consistent with the findings from Figure 1, FedProx with smaller $\lambda$ (greater personalization) consistently achieves faster convergence at the early stages of training, while FedProx with larger $\lambda$ (less personalization) takes more communication rounds to converge. As statistical heterogeneity increases (measured by $R$), FedProx with smaller $\lambda$ gradually outperforms models with larger $\lambda$, showcasing the adaptability of FedProx to different levels of heterogeneity. These results further validate the theoretical insights from Theorem 1 and Theorem 2, highlighting the influence of personalization on statistical and optimization performance in federated setups with larger numbers of clients.
Figure 6: Effect of local updates on FedProx performance for local models in MNIST logistic regression and global models ni MNIST CNN, complementing the results shown in Figure 2.
Figure 6 demonstrates the impact of varying the number of local updates on FedProx performance, complementing the results shown in Figure 2. For FedProx with smaller $\lambda$ (higher personalization), increasing the number of local updates notably reduces the need for communication, leading to faster convergence. In contrast, for FedProx with larger $\lambda$ (lower personalization), increasing local updates yields minimal benefit and only adds to computation costs without reducing communication demands. The results indicate that for larger $\lambda$ , the optimal strategy is to prioritize frequent communication rather than increasing local computation. This highlights the trade-off between local computation and communication efficiency under different levels of personalization, validating our theoretical findings in Corollary 1.
Dataset Client Class One-stage FedProx GlobalTrain LocalTrain pFedMe
(small) (medium) (large) (small) (medium) (large)
@@ C.2 Additional Experiment Results
To further validate our conclusions about the effect of personalization on the statistical accuracy of FedProx (Objective 1), we conduct additional experiments on real datasets, including EMNIST and CIFAR-10. For each dataset, we explore three levels of statistical heterogeneity, controlled by varying the number of classes each client has access to, which follows a widely used setup in federated learning literature. The models are trained using different convolutional neural network (CNN) architectures, denoted as CNN2 and CNN5, where the numbers indicate the number of convolutional layers in the network. As the true underlying model is not accessible for these real datasets, we evaluate model performance based on classification accuracy on the test set, as shown in Tables 3 and 4. We also conduct experiments another regularization-based personalization method, pFedMe (T Dinh et al., 2020), to demonstrate that our results could be generalized to other regularization-based personalized federated learning methods as well. For the columns corresponding to One-stage FedProx and pFedMe, we conduct experiments with varying regularization parameter $\lambda$, investigating the role of personalization in the algorithms. Specifically, we experiment with small, medium, and large $\lambda$ values (see Appendix C.1 for detailed parameter settings).
@@
In Table 3, we can see that consistently across datasets, as $\lambda$ decreases (indicating a higher degree of personalization), the accuracy approaches that of LocalTrain, which trains separate models for each client. Conversely, as $\lambda$ increases (indicating lower personalization and more collaboration), the accuracy approaches that of GlobalTrain, which represents purely global training. This interpolation behavior between LocalTrain and GlobalTrain, enabled by FedProx’s personalized objective, aligns well with our theoretical results in Theorem 1 and highlights the adaptive capability of the FedProx method.
@@
Moreover, as dataset homogeneity increases (i.e., clients share more similar data distributions), we observe a general improvement in accuracy across all methods, regardless of the level of personalization. This result demonstrates that while personalization offers significant benefits in highly heterogeneous settings, its necessity diminishes as heterogeneity decreases. Additionally, experiments with different CNN architectures confirm the robustness of these findings across model designs. Although we did not provide a theoretical analysis for pFedMe in this paper, our results suggest that the empirical behavior of pFedMe is consistent with the theoretical insights derived for FedProx. In particular, we observe that pFedMe also exhibits interpolation between local and global learning under varying levels of data heterogeneity.
Dataset Client Class One-stage FedProx (small $\lambda$) One-stage FedProx (large $\lambda$) Dichotomous Strategy
MNIST CNN2 2 0.4721 0.6019 0.2567
MNIST CNN2 6 0.7833 0.7843 0.7001
MNIST CNN2 10 0.8893 0.9032 0.8753
MNIST CNN5 2 0.8891 0.9333 0.7032
MNIST CNN5 6 0.8451 0.8941 0.7941
MNIST CNN5 10 0.9091 0.9102 0.8922
EMNIST CNN5 30 0.6084 0.6089 0.5291
EMNIST CNN5 40 0.6097 0.6123 0.5732
EMNIST CNN5 47 0.6278 0.6415 0.5856
Table 4: Comparison of evaluation accuracy for One-stage FedProx with different values of $\lambda$ and Dichotomous Strategy across various datasets and client classes.
C.3 Comparison of FedProx with A Dichotomous Strategy
A simple baseline in personalized federated learning is the dichotomous strategy, where the dataset on each client is split into two subsets: one is used for global training (GlobalTrain) with pooled client data, while the other is used for purely local training (LocalTrain). After training, the two models are compared on the test set, and the best-performing model is selected for each client. Such a strategy is known to achieve minimax-optimal statistical accuracy under certain conditions; see Chen et al. (2023b) for a detailed discussion.
Even though the convergence rate coincides with such a dichotomous strategy, we may reiterate that the key contribution of our work lies in the fine-grained analysis of the FedProx method—a method that has demonstrated strong empirical performance in federated learning [2]. Through our analysis, we have, for the first time, established the minimax optimality of FedProx, showing that its statistical performance matches the lower bound of the estimation problem. If one seeks to understand why FedProx could be advantageous over the dichotomous strategy in practice, here are some of the reasons:
  • Stability in Heterogeneous Environments: Empirical studies [1] demonstrate that FedProx achieves more stable convergence compared to methods such as FedAvg, especially in settings with significant in a heterogeneous network.
  • Adaptivity and Flexibility: FedProx seamlessly interpolates between the extremes of global training and local training by tuning a single regularization parameter, $\lambda$. This adaptability allows practitioners to dynamically balance the trade-off between bias and variance based on the level of statistical heterogeneity, without the need for fixed, hard decisions inherent in the dichotomous strategy.
To further compare FedProx with the dichotomous strategy, we conduct a simple experiment. In the dichotomous setup, each client’s dataset is partitioned into two subsets: one subset is used for local training (LocalTrain), where each client trains independently, and the other subset is pooled across all clients for global training (GlobalTrain). The accuracy of the two resulting models is evaluated on a test set, and the better-performing model is reported for each client. By contrast, FedProx is implemented using the experimental setup described earlier, leveraging the full dataset. As shown in Table 4, FedProx consistently outperforms the dichotomous strategy across various configurations. This performance gap can be attributed to the inefficiency of the dichotomous approach, which sacrifices valuable data during the sample splitting process, resulting in reduced sample efficiency.
Even in a hypothetical scenario where the full sample is used for both LocalTrain and GlobalTrain in the dichotomous strategy, the results in Table 3 demonstrate that FedProx achieves a continuous interpolation between the two extremes of LocalTrain and GlobalTrain by adjusting its regularization parameter $\lambda$. This aligns with our theoretical findings.

### v4

@@ D.1 More Experiment Details
Evaluation Criterion. In terms of evaluation, for the synthetic dataset, we track the ground truth local model ${{\bm{w}}_{\star}^{(i)}}$. We compute the local statistical error as $\|\widetilde{\bm{w}}^{(i)}-{{\bm{w}}_{\star}^{(i)}}\|^{2}$. To capture the total error at the $t$-th iteration, we report the local total error $\|{\bm{w}}_{t,0}^{(i)}-{{\bm{w}}_{\star}^{(i)}}\|^{2}$, which combine both optimization and statistical errors. For the MNIST dataset, we report test accuracy on the MNIST testing set as a measure of the statistical error, and training loss for each client’s training set as total error at the $t$-th iteration.
D.2 Additional Experiment Results
The Effect of Personalization on Accuracy over different data heterogeneity To further validate our conclusions about the effect of personalization on solving Problem 2, we conduct additional experiments on real datasets, including EMNIST and CIFAR-10. For each dataset, we explore three levels of statistical heterogeneity, controlled by varying the number of classes each client has access to, which follows a widely used setup in federated learning literature. The models are trained using different convolutional neural network (CNN) architectures, denoted as CNN2 and CNN5, where the numbers indicate the number of convolutional layers in the network. As the ground truth underlying model is not accessible for these real datasets, we evaluate model performance based on classification accuracy on the test set, as shown in Tables 4. We also conduct experiments another regularization-based personalization method focusing on Problem (2), pFedMe (T Dinh et al., 2020), to demonstrate that our results could be generalized to other regularization-based personalized federated learning methods as well. For the columns corresponding to FedCLUP and pFedMe, we conduct experiments with varying regularization parameter $\lambda$, investigating the role of personalization in the algorithms. Specifically, we experiment with small, medium, and large $\lambda$ values (see Appendix D.1 for detailed parameter settings).
Dataset Client Class FedCLUP GlobalTrain LocalTrain pFedMe
high median low high median low
MNIST Logit 2 0.5001 0.6011 0.6109 0.7001 0.3214 0.5534 0.6312 0.6712
6 0.6813 0.7001 0.7241 0.7718 0.6000 0.6312 0.7123 0.7681
10 0.8123 0.8213 0.8312 0.8391 0.7828 0.8001 0.8239 0.8291
@@ D.2 Additional Experiment Results
In Table 4, we can see that consistently across datasets, as personalization degree increase, the accuracy approaches that of LocalTrain, which trains separate models for each client. Conversely, as personalization degree decreases, the accuracy approaches that of GlobalTrain, which represents purely global training. This interpolation behavior between LocalTrain and GlobalTrain, enabled by Problem (2), aligns well with our theoretical results in Theorem 1.
@@
Moreover, as dataset homogeneity increases (i.e., clients share more similar data distributions), we observe a general improvement in accuracy across all methods, regardless of the level of personalization. This result demonstrates that while personalization offers significant benefits in highly heterogeneous settings, its necessity diminishes as heterogeneity decreases. Additionally, experiments with different CNN architectures confirm the robustness of these findings across model designs. Although we did not provide a theoretical analysis for pFedMe in this paper, our results suggest that the empirical behavior of pFedMe is consistent with the theoretical insights derived for FedCLUP. In particular, we observe that pFedMe also exhibits interpolation between local and global learning under varying levels of data heterogeneity. We observe in Table 4 that LocalTrain consistently underperforms across all levels of statistical heterogeneity. This is because, while statistical heterogeneity is introduced by restricting each client’s exposure to only a subset of classes in the training process, evaluation is conducted on the entire dataset, including classes not seen during training. As a result, LocalTrain’s performance remains lower than that of other methods.
The Effect of Personalization on Statistical Error over different data heterogeneity Figure 4 demonstrates the effect of personalization on statistical error across varying levels of statistical heterogeneity $R$. As demonstrated in the theory, increasing personalization (moving from FedCLUP (low per) to FedCLUP (high per)) shifts the solution closer to LocalTrain, while decreasing personalization makes the model behave more like GlobalTrain. For low heterogeneity, methods with lower personalization (i.e., greater collaboration) achieve lower statistical error, as information can be effectively shared across clients to improve learning. GlobalTrain achieves the lowest error in this regime, while FedCLUP with less personalization also performs well. However, as statistical heterogeneity increases, collaborative learning becomes less effective. In this setting, LocalTrain outperforms other methods. Highly personalized FedCLUP also performs better than less personalized variants, aligning with the theory we established in Theorem 1.
Figure 4: Impact of personalization on statistical error of FedCLUP solution across different levels of heterogeneity (high per means training FedCLUP with high personalization degree).
In Figure 5, we track testing loss as a function of communication rounds. Across all sub-figures, FedCLUP with smaller $\lambda$ (higher personalization) demonstrates faster convergence, especially in the early stages of training. Thus, it shows that the trade-off between statistical accuracy and computation cost under the influence of personalization widely exists over statistical heterogeneity.
Figure 5: Logistic Regression training loss versus communication rounds for the FedCLUP methods on MNIST dataset (high per refers to training FedCLUP with high personalization degree).
Figure 6: Performance of FedCLUP on the Synthetic dataset with 100 clients (top row) and 1,000 clients (bottom row) with 10% client sampling under different levels of statistical heterogeneity ($R$).
Different Number of Clients In Figure 6, the top row shows results for 100 clients, and the bottom row shows results for 1,000 clients, and in both results we adopt a client sampling of 10% in each communication round. This tests the scalability and generalizability of the our results under typical federated learning setups. Consistent with the findings from Figure 1, FedCLUP with smaller $\lambda$ (higher personalization) consistently achieves faster convergence at the early stages of training, while FedCLUP with larger $\lambda$ (lower personalization) takes more communication rounds to converge. As statistical heterogeneity increases (measured by $R$), FedCLUP with smaller $\lambda$ gradually outperforms models with larger $\lambda$, showcasing the adaptability of solving Problem 2 to different levels of heterogeneity. These results further validate the theoretical insights from Theorem 1 and Theorem 2, highlighting the influence of personalization on statistical and optimization performance in federated setups with larger numbers of clients.
The Effect of Personalization on Communication Cost Figure 2 investigates how personalization impacts communication efficiency in FedCLUP by analyzing the benefit of increasing local updates under different personalization levels. In the low-personalization setting (left column), more local updates significantly accelerate convergence, reducing the reliance on frequent communication. However, in the high-personalization setting (right column), increasing local updates has a limited effect on convergence, indicating that frequent communication is essential for effective learning. This demonstrates that higher personalization requires more communication rounds to achieve comparable performance, aligning with our theoretical findings.
D.3 Additional Experiment for Rebuttal
D.3.1 Strongly Convex Setting with Synthetic Data
As the theoretical results is established under strongly-convex settings, hence some expiremnts under strongly convex setting is expected with discussion related to the conditional number $\kappa$. We conduct an experiment on strongly convex problem: an overdeterminded linear regression task. We strictly follow the choice of local step size, local computation rounds, and global step size as specified in Corollary 2. As shown in Table 7 and 8, for a fixed personalization parameter $\lambda$, we observe that longer value of $\kappa$ result in slower convergence rates with respect tot he number of communication rounds. This expirical trend is consistent with our theoretical prediction in Corollary 2, where the number of communication rounds required to achieve a given target error $\epsilon$ scales with $\mathcal{O}(\kappa\frac{\lambda+L}{\lambda+\mu}\log(1/\varepsilon))$. Moreover, we observe that the impact of increasing $\kappa$ becomes stronger as $\lambda$ increases. This phenomenon aligns with our theoretical analysis, which shows that the sensitivity of communication complexity to $\kappa$ is amplified for larger values of $\lambda$. Similarly, we observe that increasing the condition number $\kappa$ also leads to a higher total number of gradient evaluations required to reach a target error. This observation is in agreement with our theoretical results where the total number of local gradient evaluations scales as $\mathcal{O}(\kappa\log(1/\varepsilon))$.
Figure 7: Optimization error versus communication rounds under varying condition numbers $\kappa$. The left, middle, and right panels correspond to FedCLUP with small, medium, and large values of $\lambda$, respectively. Across all settings, we observe that larger $\kappa$ leads to slower convergence, with the effect becoming more pronounced as $\lambda$ increases.
Figure 8: Optimization error versus total number of local gradient evaluations per client under varying condition numbers $\kappa$. As $\kappa$ increases, the number of gradient evaluations required to achieve a given optimization error also increases.
D.3.2 Datasets with Natural Partitions
(a) The loss of FedCLUP with CelebA
(b) The loss of pFedMe with CelebA
Figure 9: Loss of FedCLUP (a) and pFedMe (b) with different personalization degrees on the CelebA dataset. Low, median, and high personalization correspond to regularization terms 0.5, 0.1, and 0.001, respectively. For FedCLUP, the test accuracy is highest with low personalization (0.919), followed by median (0.915) and high (0.888). For pFedMe, the test accuracy is highest for low personalization (0.910), followed by median (0.901) and high (0.890). The test accuracy for GlobalTrain and LocalTrain is 0.911 and 0.561, respectively.
(a) The loss of FedCLUP with Sent140
(b) The loss of pFedMe with Sent140
Figure 10: Loss of FedCLUP (a) and pFedMe (b) with different personalization degrees on the Sent140 dataset. Low, median, and high personalization correspond to regularization terms 0.1, 0.005, and 0.001, respectively. For FedCLUP, the test accuracy is highest for low personalization (0.734), followed by median (0.702) and high (0.696). For pFedMe, the test accuracy is highest for low personalization (0.727), followed by median (0.703) and high (0.701). The test accuracy for GlobalTrain and LocalTrain is 0.730 and 0.353, respectively.